OpenStudy (anonymous):

Simplify. 25/55x 1.5/11x 2.1/2x 3.x/2 4.5x/12

OpenStudy (anonymous):

5/11x

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

Is itthe problem \(\dfrac{25}{55x} \) or \(\dfrac{25}{55}x\) ?

OpenStudy (anonymous):

You need to understand why.

OpenStudy (anonymous):

Whatever

OpenStudy (anonymous):

yah the first one math student

OpenStudy (mathstudent55):

\(\dfrac{25}{55x} = \dfrac{5 \times 5}{5 \times 11x} = \dfrac{5}{5} \times \dfrac{5}{11x} =\) \(= 1 \times \dfrac{5}{11x} = \dfrac{5}{11x}\)

OpenStudy (mathstudent55):

You can reduce the fraction by dividing the numerator and the denominator by 5 as shown above.

OpenStudy (anonymous):

ooh thanks :)

OpenStudy (mathstudent55):

yw